Please help me solve these:
Are the roots given for each equation correct?
w squared + 14w = 49 (7, -7) (I say this is not correct but I can't prove why)
2t squared - 7t -4 = 0 (4, - 1/2) ____________________(I don't know how to solve this)

Answer by stanbon(75887)

Are the roots given for each equation correct?
Substitute the number pair into the equation to see if you get a true result.
----------------------------------------------------------------------------
w squared + 14w = 49 (7, -7)
w^2 + 14w = 49
7^2 + 14*7 = 49
49 + 98 = 49
147 = 49
No, so 7 is not a root.
-------------------------------
(-7)^2 +14*-7 = 49
49 - 98 = 49
-50 = 49
No, so -7 is not a root.
------------------------------------
2t squared - 7t -4 = 0 (4, - 1/2)
2t^2 - 7t - 4 = 0
2(4)^2 - 7*4 - 4 = 0
32 - 28 - 4 = 0
0 = 0
Yes, so 4 is a root.
----------------------------
2(-1/2)^2 - 7(1/2) - 4 = 0
2(1/4) - (7/2) - 4 = 0
1/2 - 3 1/2 - 4 = 0
No, so -1/2 is not a root
===========================
Cheers,
Stan H.